use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions
s2 - 4s + 4 = 0

Answer by richard1234(7193)

Discriminant = b^2 - 4ac = (-4)^2 - 4(1)(4) = 0.

There is one (real) double root.